### v2

# for Large Language Models

Abstract
Previous multilingual benchmarks focus primarily on simple understanding tasks, but for large language models (LLMs), we emphasize proficiency in instruction following, reasoning, long context understanding, code generation, and so on. However, measuring these advanced capabilities across languages is underexplored. To address the disparity, we introduce BenchMAX, a multi-way multilingual evaluation benchmark that allows for fair comparisons of these important abilities across languages. To maintain high quality, three distinct native-speaking annotators independently annotate each sample within all tasks after the data was machine-translated from English into 16 other languages. Additionally, we present a novel translation challenge stemming from dataset construction. Extensive experiments on BenchMAX reveal varying effectiveness of core capabilities across languages, highlighting performance gaps that cannot be bridged by simply scaling up model size. BenchMAX serves as a comprehensive multilingual evaluation platform, providing a promising test bed to promote the development of multilingual language models. The dataset [^1] https://huggingface.co/collections/LLaMAX/benchmax-674d7a815a57baf97b5539f4 and code [^2] https://github.com/CONE-MT/BenchMAX.git are publicly accessible.
BenchMAX: A Comprehensive Multilingual Evaluation Suite
for Large Language Models
Xu Huang1, Wenhao Zhu1, Hanxu Hu2, Conghui He3, Lei Li4, Shujian Huang1thanks: Corresponding authors, Fei Yuan311footnotemark: 1 1National Key Laboratory for Novel Software Technology, Nanjing University 2University of Zurich,4Carnegie Mellon University 3Shanghai Artificial Intelligence Laboratory {xuhuang,zhuwh}@smail.nju.edu.cn, hanxu.hu@uzh.ch, heconghui@pjlab.org.cn leili@cs.cmu.edu, huangsj@nju.edu.cn, yuanfei@pjlab.org.cn
1 Introduction
Large Language Models (LLMs;  OpenAI et al., 2024; Gemini, 2024; DeepSeek-AI et al., 2024) have displayed remarkable proficiency across a wide range of tasks, mainly because they excel in instruction following, reasoning, long context understanding, code generation, and so on Ouyang et al. (2022); Cobbe et al. (2021); Su et al. (2024); Roziere et al. (2023); Lu et al. (2024); Sun et al. (2024). Inherently, these capabilities are language-agnostic. Consider a simple task like the acquisition of mathematical concept: the numerical outcome remains consistent regardless of whether one learns the arithmetic expression $1+1=2$ in English or Chinese. Similarly, when it comes to coding tasks, the choice between English or Chinese for articulating these instructions does not alter the fundamental logic of the code. However, numerous empirical studies have shown that LLMs’ multilingual performance is quite unbalanced across different languages when handling same tasks Shi et al. (2023); Zhu et al. (2024); Qi et al. (2023).
Figure 1: BenchMAX evaluates diverse advanced capabilities of LLMs in multilingual context.
However, current benchmarks Hendrycks et al. (2021); Lai et al. (2023); Singh et al. (2024); Wang et al. (2024a) do not support comprehensive testing of the language-agnostic abilities of LLMs, particularly in low-resource language settings, for several reasons. Tasks like XWinograd Muennighoff et al. (2023) and XStoryCloze Lin et al. (2022), based on multiple-choice formats, do not fully evaluate the generative capacities of LLMs. Additionally, the limited language overlap across existing benchmarks poses challenges in assessing LLM performance in diverse languages. Recently, P-MMEval Zhang et al. (2024) is proposed as a multilingual multitask benchmark, with the majority of its tasks still following a multiple-choice format. While it includes assessments like MGSM Shi et al. (2023) and MIFEVAL that cover partial language-agnostic capabilities, LLMs exhibit remarkable performance, as shown in Table 1. This narrow focus leaves a significant gap between research evaluation and real-world applications.
To tackle this problem, we develop a comprehensive, multi-way, and challenging multilingual evaluation suite, called BenchMAX, to help the community better analyze and improve the language-agnostic capabilities of LLMs. Covering 17 languages [^3] The 17 languages include English, Spanish, French, German, Russian, Bengali, Japanese, Thai, Swahili, Chinese, Telugu, Arabic, Korean, Serbian, Czech, Hungarian, and Vietnamese., BenchMAX not only includes a broader range of language families but also emphasizes the diversity of writing systems across languages. As demonstrated in Table 1, BenchMAX increases the percentage of studied languages that utilize the non-Latin script.
Tasks Llama3.1 70B Qwen2.5 72B # LG # LG-Family Diversity $\boldsymbol{R_{\mathrm{non-Latin}}}$
XWinograd 69.7 83.7 6 3 50.0
XStoryCloze 70.3 83.6 13 11 38.5
MGSM 88.3 79.2 10 7 50.0
MIFEVAL 91.0 87.6 10 7 50.0
BenchMAX
- Instruction Following 11.1 34.1 17 11 58.8
- Code Generation 29.8 45.5
- Science Reasoning 35.8 39.4
- Tool Use 44.3 61.8
Table 1: BenchMAX provides a more comprehensive analysis of LLM language-agnostic capabilities by covering a broader range of capability scenarios, language families, and script systems. # LG and # LG-Family denote the number of supported languages and the language families they belong to, respectively. $R_{\mathrm{non-Latin}}$ refers to the proportion of languages that do not use the Latin script among supported languages. * The results are from P-MMEval.
Meanwhile, BenchMAX highlights diverse language-agnostic advanced capabilities (Figure 1). We assess instruction following capability with rule-based Zhou et al. (2023) and model-based Li et al. (2024) evaluations, code generation capability in diverse scenarios (function-completion Liu et al. (2024) / problem solving Jain et al. (2024)), long context understanding capability Hsieh et al. (2024), a verity of reasoning in math Shi et al. (2023) and science Rein et al. (2023), tool use Srinivasan et al. (2023) in agent environments, and general Costa-jussà et al. (2022) / domain translation. Domain translation, a by-product of data construction, poses a new challenge for LLM by necessitating fine-grained control and domain-specific terminology understanding over the translation process.
To ensure high quality, we devise an annotation framework to optimize the dataset quality with human effort and LLM feedback. The process involves translating data from English to selected non-English languages using machine translation systems, post-editing each sample by three native-speaking annotators with multiple iterations across most tasks, and picking the final translation version using a strong LLM that involves swapping sample positions for debiasing Wang et al. (2024b); Li et al. (2024).
Leading multilingual LLMs are evaluated on BenchMAX, revealing that language significantly influences language-agnostic capabilities of existing LLMs. Interestingly, simply increasing the parameters can boost average performance on these tasks but does not universally reduce the performance gap across languages. Moreover, compared to general translation, domain translation not only poses new challenges for LLMs but also requires new evaluation metrics. The main contributions can be summarized as follows:
  • We develop a comprehensive, multi-way multilingual benchmark across 17 languages for evaluating 6 crucial capabilities on 10 diverse tasks.
  • We propose a pipeline for curating high-quality mutlilingual datasets, involving both human annotation and LLM-as-a-judge.
  • We evaluate leading multilingual LLMs on BenchMAX, and the related analyses provide a further understanding of the language-agnostic capabilities.
Language ISO Language Family Script System Language ISO Language Family Script System
Hungarian hu Uralic Latin Serbian sr Indo-European Serbian Cyrillic
Vietnamese vi Austroasiatic Korean ko Koreanic Hangul / Chosŏn’gŭl
Spanish es Indo-European Japanese ja Japonic Mixed scripts of Chinese Characters and Hiragana, Katakana
Czech cs Arabic ar Afro-Asiatic Arabic alphabet
French fr Thai th Kra–Dai Thai
German de Swahili sw Niger–Congo Latin
Russian ru Cyrillic Chinese zh Sino-Tibetan Chinese Characters
Bengali bn Bengali–Assamese Telugu te Dravidian Telugu
Table 2: Besides English, BenchMAX supports 16 non-English languages, covering a wide range of language families and script systems.
Capability Category Dataset # Samples Metric Capability Category Dataset # Samples Metric
Instruction Rule-based IFEval 429 Accuracy Code Generation Function Completion Humaneval+ 164 Pass@1
Following Model-based m-ArenaHard 500 Win Rate Problem Solving LiveCodeBench_v4 713
Reasoning Math MGSM 250 Exact Match Translation General Flores+TED+WMT24 [[[[1012, 4049]]]] spBLEU
Science GPQA 448 Domain Annotated data above 2781
Tool Use Multiple Functions Nexus 318 Accuracy Long Context Modeling Question Answering RULER 800 Exact Match
Table 3: Selection of core capabilities and details of task data. For IFEval, we filter out all language-specific instructions, thus remaining 429 samples. For Nexus, we only adopt the standardized_queries subset which contains 318 samples. For general translation datasets, the number of samples may vary in different translation directions, according to the number of parallel samples in TED and WMT24. The datasets of the model-based instruction following task and math reasoning are expanded from existing multilingual datasets, while others are translated from English datasets.
2 Related Work
Prior to the era of Large Language Models  (LLMs;  OpenAI et al., 2024; Gemini, 2024; DeepSeek-AI et al., 2024), most multilingual benchmarks are designed to evaluate discriminative models and take the form of classification tasks, such as XNLI Conneau et al. (2018), XCOPA Ponti et al. (2020), XCSQA Talmor et al. (2019), and so on Lin et al. (2022); Muennighoff et al. (2023). However, due to their limited task complexity and the lack of diversity in format, these tasks become less practical. Recently, MGSM has become the most frequently used dataset in papers and reports from leading LLM teams Dubey et al. (2024); Gemini (2024); OpenAI (2024), which measures the mathematical reasoning capability across eleven languages. In this paper, we extend it to cover six additional diverse languages. Another widely used benchmark is the multilingual translated version of MMLU Hendrycks et al. (2021); Lai et al. (2023); Singh et al. (2024), which assesses LLMs on knowledge-intensive tasks. However, due to the lack of a unified dataset version, scores are often difficult to compare across studies. Moreover, recent analyses have revealed that MMLU contains numerous ground truth errors Gema et al. (2024), obscuring the accurate evaluation. More recently, Include Romanou et al. (2024) is proposed to evaluate the multilingual regional knowledge, hence it’s questions are not language-agnostic. To address these limitations, our work builds upon GPQA dataset Rein et al. (2023) instead of MMLU, which offers higher-quality annotations and poses greater challenges in domain-specific knowledge and reasoning evaluation. In addition to curating multilingual versions of MGSM and GPQA, we incorporate a broader range of capabilities, including long context modeling Hsieh et al. (2024), tool use Srinivasan et al. (2023), instruction following Zhou et al. (2023), and more. This allows our benchmark to evaluate LLMs’ multilingual capabilities more comprehensively compared to previous aggregated benchmarks, such as SeaEval Wang et al. (2024a) and P-MMEval Zhang et al. (2024). More importantly, all translations except the long context data in our benchmark are post-edited by native human experts after machine translation. This significantly improves both the quality and reliability of the dataset.
3 Benchmark Construction
In this section, we extend the evaluation of the core capabilities of LLMs into multilingual scenarios. To ensure sufficient linguistic diversity, we select 16 non-English languages (§ 3.1). Meanwhile, a diverse set of tasks designed to evaluate 6 crucial LLM capabilities is chosen to facilitate comprehensive assessment (§ 3.2). Subsequently, we introduce a rigorous pipeline (§ 3.3) that incorporates human annotators and LLMs to obtain a high-quality benchmark.
Figure 2: The construction process involves three steps: Step 1) translating data from English to non-English; Step 2) post-editing each sample by three human annotators; Step 3) selecting the final translation version.
3.1 Language Selection
BenchMAX supports 17 selected languages to represent diverse language families and writing systems (Table 2).
Figure 3: Flow chart illustrating the constraint extraction and machine translation pipeline in the first step of our benchmark construction.
3.2 Capabilities Selection
LLMs have demonstrated proficiency in understanding tasks such as text classification and sentiment analysis, but their capabilities transcend understanding. We construct tasks to evaluate following intrinsic capabilities in multilingual settings:
  • Instruction Following: The capability to follow instructions is evaluated by two distinct tasks with different evaluation paradigms: rule-based and model-based assessment. For the rule-based task, we translate IFEval Zhou et al. (2023) from English to other languages, while we expand m-ArenaHard Dang et al. (2024); Li et al. (2024) to languages we selected for the model-based task.
  • Reasoning: The reasoning capability is assessed through intricate scenarios including math and natural science (physics, chemistry, and biology) problems. We expand MGSM Shi et al. (2023) and GPQA Rein et al. (2023) to 17 languages for the math reasoning and science reasoning tasks.
  • Code Generation: We primarily consider Python code generation in two settings, function completion and programming problem solving. We translate Humaneval+ Liu et al. (2024); Chen et al. (2021) and LiveCodeBench_v4 Jain et al. (2024) from English to other languages.
  • Long Context Modeling: We evaluate the ability to extract evidence from lengthy documents through question-answering tasks with long documents (128k tokens). We build this task based on RULER Hsieh et al. (2024), and translate haystacks, needles, and QA pairs.
  • Tool Use: We assess the ability to correctly select and invoke a single function from multiple available options in response to user queries. We translate the queries in Nexus Srinivasan et al. (2023) to other languages, but leave the APIs in English.
  • Translation: Translation involves accurately converting text between languages while preserving semantic meanings. Beyond traditional translation tasks including Flores, TED, and WMT Costa-jussà et al. (2022); Cettolo et al. (2012); Kocmi et al. (2024), we introduce the Domain Translation task, a by-product of the BenchMAX construction process. This task challenges models to translate specialized terminology and determine whether specific segments should be translated.
The information of the seleted datasets, sample sizes, and evaluation metrics is provided in Table 3. More details can be found in Appendix A.
[Original Text] {prompt: Create an ad copy by expanding “Get 40 miles per gallon on the highway” in the form of a QA with a weird style. Your response should contain less than 8 sentences. Do not include keywords ‘mileage’ or ‘fuel’ in your response.
instruction_id_list: [‘length_constraints: number_sentences’, ‘keywords: forbidden_words’]
kwargs: [{’relation’: ‘less than’, ‘num_sentences’: 8}, {‘forbidden_words’: [‘mileage’, ‘fuel’]}]}
[Translation Input] Create an ad copy by expanding "Get 40 miles per gallon on the highway" in the form of a QA with a weird style. Your response should contain less than 8 sentences. Do not include keywords ‘$<$b$>$mileage$<$/b$>$’ or ‘$<$b$>$fuel$<$/b$>$’ in your response.
[Google Translation Result] {CJK}UTF8gbsn以风格怪异的问答形式扩展“在高速公路上每加仑行驶 40 英里”来创建广告文案。您的回复应少于 8 个句子。请勿在回复中包含关键字“<<<b>>>里程<<</b>>>”或“<<<b>>>燃料<<</b>>>”。
[Postprocessing] {prompt: {CJK}UTF8gbsn以风格怪异的问答形式扩展“在高速公路上每加仑行驶 40 英里”来创建广告文案。您的回复应少于 8 个句子。请勿在回复中包含关键字“里程”或“燃料”。
instruction_id_list: [‘length_constraints: number_sentences’, ‘keywords: forbidden_words’]
kwargs: [{‘relation’: ‘less than’, ‘num_sentences’: 8}, ‘forbidden_words’: [{CJK}UTF8gbsn‘里程’, ‘燃料’]] }
[Human Post-Editing] {prompt: {CJK}UTF8gbsn以一种奇特风格的问答形式展开“在高速公路上每加仑行驶40英里”这句话,创建为一个广告文案。你的回答应该少于8句话。不要在你的回复中包含关键字“里程”或“燃料”。
instruction_id_list: [‘length_constraints: number_sentences’, ‘keywords: forbidden_words’]
kwargs: [{‘relation’: ‘less than’, ‘num_sentences’: 8}, ‘forbidden_words’: [{CJK}UTF8gbsn‘里程’, ‘燃料’]] }
Table 4: One example in rule-based instruction following task, which includes complex constraints. First, we enclose these constraints with special symbols and then translate the prompt from English to the target language by Google Translate. Finally, we postprocess the prompt by extracting constraints into kwargs and removing special symbols for human post-editing.
Setting Target Language
zh es fr hu
w/o special symbols 0.68 0.68 0.68 0.68
symbol 1: $<$b$>$ $<$/b$>$ 0.91 0.89 0.88 0.93
symbol 2: ( ) 0.88 0.91 0.89 0.92
symbol 3: ([ ]) 0.82 0.89 0.87 0.92
Order 1 + symbol 1 0.91 0.89 0.88 0.93
+ symbol 2 0.93 0.93 0.90 0.95
Order 2 + symbol 2 0.88 0.91 0.89 0.92
+ symbol 1 0.90 0.93 0.90 0.95
+ symbol 3 0.92 0.93 0.90 0.95
Table 5: The recall rates of constraints using different groups of special symbols. We choose Order 1, which has fewer steps and produces on-par or better performance than other settings.
3.3 Construction
The way to obtain BenchMAX consists of three steps, as shown in Figure 2: 1) translate data from English to non-English by machines; 2) post-edit each sample by three native annotators; 3) pick the final translation version by GPT-4o-mini.
Step 1: Translating data from English to selected non-English languages by machine translation systems.
We select between traditional translators such as Google Translate, and LLM-based ones like GPT-4o, depending on whether a task contains extractable constraints. As illustrated in Figure 3, if the data contains constraints that are hard to extract, we prompt GPT-4o to translate the data and satisfy the constraints. Otherwise, we use Google Translate along with extraction tools. Extraction tools include methods for extracting translated keywords by enclosing source keywords with special symbols, and for preserving source constraints by replacing constraints with placeholders before translation and restoring them afterwards.
Taking an example of rule-based instruction following task as an example, as shown in Table 3.2, it requires extra processing to extract constraints from the translated instruction, as they are needed for verification. Inspired by Yuan et al. (2020), we enclose the keywords in the original instruction with special symbols, making them easy to extract from the translated result. If one symbol fails, another symbol is used to improve recall. As shown in Table 5, we explore various groups of special symbols and different orders, and calculate the recall rates of keywords. Comparing to not using special symbols, apply any symbol group can greatly improve the recalls, while combining different symbol groups in multiple rounds can further improve the recalls. We choose Order 1 as it can achieve better results with fewer groups than Order 2. The detailed recall results are in Appendix B.
Note that in cases where existing multilingual datasets are available, such as MGSM and m-ArenaHard, we extend them to include the supported languages by translating the English data, to minimize additional effort.
Model Size Instruction Following Code Generation Reasoning Long Context Tool Use Translation
Rule-based Model-based Func Compl. Prob. Solving Math Science Question Answering Multi Func. General Domain
En-X X-En En-X X-En
InternLM2.5 7B 45.7 1.9 45.4 10.3 37.4 20.6 37.5 53.2 12.7 20.2 34.4 54.0
20B 51.9 3.3 51.2 14.4 42.9 24.0 - 26.6 14.9 19.7 34.9 53.9
Aya-Expanse 8B 51.2 6.4 33.8 7.8 50.8 26.2 - 41.1 21.5 26.8 45.6 51.6
32B 61.9 12.4 52.0 15.8 66.7 27.7 - 59.8 25.2 32.8 54.8 62.3
Gemma2 9B 63.0 9.8 53.9 16.6 72.0 23.9 - 61.4 27.2 33.2 57.5 61.9
27B 62.4 18.0 66.7 24.6 75.3 26.7 - 64.7 30.4 34.5 64.8 66.2
Llama3.1 8B 62.6 4.3 52.9 14.1 63.4 23.8 68.3 45.0 24.6 29.8 53.9 62.9
R1-Distill-Llama3.1 8B 49.7 3.5 62.8 23.8 46.9 28.1 - 37.2 12.2 20.8 13.5 23.1
Llama3.1 70B 76.2 13.2 69.7 29.8 79.7 35.8 57.4 44.3 31.1 35.1 64.5 68.2
Llama3.3 70B 85.2 17.0 74.0 34.7 83.8 42.6 50.4 42.5 31.5 33.6 63.5 65.0
R1-Distill-Llama3.3 70B 78.0 26.6 84.6 54.8 82.8 46.1 - 62.1 26.0 33.0 47.6 45.2
Qwen2.5 7B 65.9 8.5 68.2 24.7 63.4 27.6 53.5 48.9 16.6 25.6 46.4 60.0
R1-Distill-Qwen2.5 7B 46.7 3.0 69.3 37.3 56.1 28.4 - 27.7 6.8 16.3 17.0 27.3
Qwen2.5 32B 78.1 17.3 75.8 42.7 77.7 37.7 79.4 66.7 22.7 30.5 54.2 65.4
R1-Distill-Qwen2.5 32B 67.3 19.2 80.6 54.4 77.3 37.0 - 60.4 20.3 28.5 37.1 37.7
Qwen2.5 72B 80.8 36.9 78.6 45.5 77.8 39.4 80.6 61.8 25.8 33.3 60.4 66.9
DeepSeek-V3* 671B 83.9 59.8 83.2 60.4 84.2 47.4 85.2 69.2 33.9 34.5 70.3 67.8
GPT-4o-mini - 79.1 21.9 78.7 37.0 76.9 34.1 82.1 70.9 30.3 33.9 67.7 67.6
Table 6: Performance comparison across models on BenchMAX tasks, averaged over 17 languages. Bold numbers indicate the best performance in each column. "Func Compl." refers to Function Completion, "Prob. Solving" to Problem Solving, and "Multi Func." to Multiple Functions scenarios where models must select and call one function from multiple options. Models without results on the long context task do not support 128K context length. * DeepSeek-V3 is a 671B MoE model, with 37B activated for each token.
Step 2: Post-editing each sample by three distinct native-speaking annotators in almost all tasks.
To ensure high-quality dataset, we implement a multi-round annotation and verification process. 1) Each sample is given to three native-speaking annotators who are proficient in English and their native language. Considering the specialized nature of datasets like Science reasoning, annotators are required to hold at least a Bachelor’s degree. 2) Two automatic verifiers - rule-based verifiers and model-based verifiers - are used to assess the quality of human annotation. Rule-based verifiers ensure the satisfaction of constraints for certain tasks, such as the rule-based instruction following task. For model-based verifiers, we utilize the GEMBA-SQM prompt Kocmi and Federmann (2023) and employ Qwen2.5-72B-Instruct, a powerful multilingual model, to estimate the quality of translations. Along with providing an overall score, the model offers detailed explanations of translation errors as feedback to annotators. Samples that do not pass the rule-based verifier or score below a predefined threshold are identified as failed, and refined in subsequent iterations. Each manually annotated dataset undergoes at least three iterations.
Step 3: Selecting the final translation version by LLMs.
Initially, we ask a fourth annotator, uninvolved in the annotation process, to choose the final version from the results revised by three individuals. Intriguingly, the selection by the fourth annotator exhibited a strong position bias, often favoring the initial annotation. This preference could be attributed to the uniformly high quality of translations, resulting in minimal discernible differences among them.
Debiasing for human annotators is costly in terms of both time and finance, because three translations encompass all permutations of six. Consequently, we employ GPT-4o-mini to select the final translation, as it is a powerful and balanced LLM across different languages. In particular, following Li et al. (2024), we adapt the LLM-Judge system instruction (see Appendix D) to suit pairwise translation evaluation. We shuffle the positions of three translations and conduct two battles to select a final version. In each battle between two translations, we perform two judgments by swapping their positions and determine one winner. The winner of the first two translations then battles against the third translation, determining the final winner.
4 Experimental Results
4.1 Evaluation Setup
Evaluated Models We focus primarily on multilingual post-trained models and evaluate both open-source and proprietary language models [^4] Unless otherwise specified, all models discussed in this paper are post-trained versions., including Llama3.1 Dubey et al. (2024), Qwen2.5 Qwen Team (2024), Gemma2 Team et al. (2024), InternLM2.5 Cai et al. (2024), Aya-Expanse Dang et al. (2024), DeepSeek-R1-Distill-Llama Guo et al. (2025), DeepSeek-R1-Distill-Qwen, DeepSeek-V3 DeepSeek-AI et al. (2024), and GPT-4o-mini OpenAI (2024). Models’ detailed descriptions are in Appendix C.
Inference Configuration We adopt greedy decoding for most tasks, except for the problem solving task, where the sampling temperature is set to $0.2$. The default chat template and system prompt of each model are applied. Detailed prompts are provided in Appendix D. For reasoning tasks, we adopt the zero-shot native chain-of-thought templates in LM-Evaluation-Harness Gao et al. (2024). For other tasks, we use the prompt templates provided in corresponding repositories [^5] https://github.com/EleutherAI/lm-evaluation-harness
https://github.com/LiveCodeBench/LiveCodeBench
https://github.com/evalplus/evalplus
https://github.com/NVIDIA/RULER
https://github.com/lmarena/arena-hard-auto
, and change the user inputs to other languages.
Figure 4: Taking two tasks as examples, models exhibit unbalanced multilingual capabilities. We show performance of several models on the science reasoning task and the domain translation task across different languages.
Figure 5: Larger models do not consistently have a smaller GAP. Each row shows proportions of tasks where the larger model achieves a smaller GAP versus where the smaller model performs better.
4.2 Multilingual Benchmark Results
Table 6 shows the overall average performance of each model on each multilingual task. More detailed results are in Appendix E.
Model scaling improves overall multilingual performance while language disparities persist.
As shown in Table 6, larger models consistently demonstrate enhanced multilingual capabilities across all domains, with few exceptions. However, the performance gap between English and non-English languages does not invariably diminish. We define GAP as the average performance gap between English and other languages: GAP=∑l≠enmax(s(en) - s(l), 0)n-1, where $s(l)$ denotes the score on the task with language $l$, and $n$ is the number of languages including English. As shown in Figure 5, when comparing models of different sizes, the proportion of larger models achieving smaller GAPs only slightly exceeds 0.5 for most model families. Gemma2-9B achieves smaller GAPs than Gemma2-27B on most tasks. These findings suggest that while scaling model size effectively improves overall multilingual performance, additional strategies may be needed to address the performance disparities across languages.
The effective utilization of language-agnostic capabilities remains challenging in multilingual contexts.
As illustrated in the left plot of Figure 4, models’ reasoning capabilities vary significantly across languages. Generally, models achieve better performance in dominant languages like English than in other low-resource languages. This disparity can be attributed to the fact that multilingual task execution depends not only on language-agnostic reasoning but also on language-specific capabilities such as comprehension and generation. Therefore, when operating in weak languages, it becomes difficult for a model to fully leverage its language-agnostic capabilities. Interestingly, we observe an unexpected pattern where certain models excel in specific non-dominant languages compared to English on some tasks. For example, Qwen2.5 demonstrates superior performance in Korean over English on the science reasoning task. We hypothesize that Qwen2.5’s training data includes a relatively high proportion of Korean content in scientific or reasoning-related domains. This counter-intuitive phenomenon warrants further investigation in future work.
Figure 6: Left: The translation performance is positively correlated with other multilingual performance. Spearman Correlations are computed between the performance on general/domain translation and the specific task. Right: Models in the same family have similar language performance pattern. We compute the Spearman Correlations between the performance of two models (Llama3.1 8B vs 70B, Qwen2.5 7B vs 72B) across different languages.
Model performance exhibits systematic bias towards high-resource languages.
As shown in Figure 4, the performance curves of most models exhibit significant fluctuations across languages. High-resource languages such as French and Chinese consistently outperform low-resource languages like Telugu, Swahili, and Bengali. For instance, while DeepSeek-V3 achieves over 50% accuracy in science reasoning tasks for English and French, its performance drops notably to below 40% for Telugu. This pattern can be partially attributed to development strategies - models like Aya-Expanse were not specifically optimized for the full range of languages in our evaluation. Unexpectedly, Gemma2 exhibits relatively balanced performance across most tasks (Figure 9), despite not being explicitly marketed as a multilingual model.
Translation capabilities exhibit a positive correlation with other evaluated capabilities.
We analyze the relationship between model’s English-to-X translation capability and other capabilities using Spearman correlation coefficients (the left panel of Figure 6). When calculating correlations between domain-specific translation performance and task performance, we exclusively use data from the corresponding domains. The analysis reveals that domain-specific translation performance generally exhibits stronger correlations with task performance compared to general translation capabilities. A notable exception is that in the rule-based instruction-following task, we observe an inverse scaling effect: larger LLMs produce lower-quality translations compared to their smaller counterparts. We find that larger LLMs are more likely to execute instructions rather than strictly perform translation, known as prompt injection.
Models within the same family exhibit consistent performance patterns across languages.
We calculate Spearman correlation coefficients to analyze the performance similarity between models of the same family (excluding R1-distilled models) across different languages for each task. As shown in the right panel of Figure 6, models within the same family show strong correlations across various tasks, with most correlation coefficients exceeding 0.7.
R1-distilled models exhibit enhanced multilingual reasoning and code generation capabilities, but some other capabilities, especially translation, are noticeably degraded.
As illustrated in Table 6, the performance of R1-Distill-Llama3.3-70B is comparable to DeepSeek-V3 in reasoning and code generation tasks, and is stronger than Llama-3.3-70B-Instruct. However, other capabilities like instruction following of 7B/8B models exhibit degradation to some extent. They tend to generate repeated tokens in the reasoning process when using non-English languages. The translation capabilities of both large and small distilled models decline dramatically. In addition to repeated generation, we also observe a frequent phenomenon of code-switching in translations.
5 Analysis
Llama3.1-70B Qwen2.5-72B
Translated by GT 4o-mini Ours GT 4o-mini Ours
Rule-based 66.9 53.5 76.2 71.5 57.2 80.8
Func Compl. 47.8 68.2 69.7 50.4 75.5 78.6
Science 33.7 35.1 35.8 36.9 37.8 39.4
Multi Func. 23.0 43.7 44.3 26.7 61.3 61.8
Table 7: Our pipeline provides a more accurate assessment of the multilingual performance, compared to naive translations by Google Translate(GT) and GPT-4o-mini(4o-mini), respectively.
5.1 Our pipeline provides a more accurate assessment of models’ performance
We naively translate a subset of tasks from English to other 16 languages by Google Translate and GPT-4o-mini, and then evaluate two models using this task data. We directly translate sources by Google Translate as it doesn’t support constraints, and use appropriate prompts with constraints to request GPT-4o-mini. The results in Table 7 show that models achieve higher scores generally on our translated tasks compared to naive machine-translated ones. Google Translate lacks the flexibility to handle diverse constraints and specific domains, while GPT-4o-mini does not always perform translation task, especially on instruction data. This indicates that naive machine translation underestimates LLMs’ capabilities, whereas our data provides a more accurate assessment.
5.2 BenchMAX is adequate to evaluate multilingual capabilities of LLMs
Prior work like Aya-Expanse Dang et al. (2024) relies on conventional understanding tasks such as XCOPA and XWinograd for multilingual evaluation. On these metrics, Gemma2-9B achieves the best performance, followed by Aya-Expanse-8B, Llama3.1-8B, and Qwen2.5-7B. However, our evaluation through BenchMAX reveals a different pattern: Qwen2.5-7B demonstrates superior multilingual capabilities on generation tasks, while Aya-Expanse models show notably weaker performance on code generation tasks, as shown in Table 8. This discrepancy highlights the importance of comprehensive evaluation frameworks that incorporate both understanding and generation tasks to accurately assess multilingual capabilities of LLMs.
Model Rule-based Func Compl. Discriminative
Qwen2.5-7B No.1 No.1 No.4
Llama3.1-8B No.3 No.3 No.3
Aya-Expanse-8B No.4 No.4 No.2
Gemma2-9B No.2 No.2 No.1
Table 8: Rankings of the models in generative tasks in BenchMAX differ from that in discriminative tasks, indicating the importance of both types of tasks. The ranking in discriminative tasks is from Dang et al. (2024).
5.3 BenchMAX reveals the challenges in domain-specific translation evaluation
Domain-specific texts often contain substantial content, such as code, that does not require translation, leading to inflated spBLEU scores. To address this limitation, we explore alternative evaluation approaches: the edit-distance metric TER Snover et al. (2006), the model-based metric xCOMET-XXL Guerreiro et al. (2024), and a novel performance retention rate that compares downstream task performance between model self-translation and human translations. Table 9 presents these metric scores across selected tasks and languages. Traditional metrics prove unreliable for domain-specific translation evaluation. Both spBLEU and TER show extreme values in Science and Programming tasks due to large portions of unchanged text, failing to capture the quality of crucial translated segments. The model-based metric xCOMET scores show inconsistency, from 18 to 96, across different scenarios, particularly struggling with low-resource languages and specialized domains. Moreover, the performance retention rate doesn’t work as expected as many values are very close, which fails to evaluate translations directly and subtly. These findings highlight the need for specialized metrics for domain-specific translation evaluation, which we identify as an important direction for future research.
Figure 7: Advanced models show high consistency between English and other languages across six tasks.
5.4 High consistency between the questions answered correctly/incorrectly in English and in other languages
Although sometimes similar performance can be achieved across different languages for certain tasks, the specific problems being addressed may vary significantly. To examine the language alignment, we compute the consistency between the problem-solving correctness in English versus other languages. Consistency is calculated as the proportion of predictions where a model’s output is correct or incorrect in both languages, out of all evaluated samples. Figure 7 presents the consistency between English and languages, based on results of Llama3.1-70B and DeepSeek-V3 on six subtasks of BenchMAX. Both these strong multilingual models demonstrate high consistency on most tasks, with most scores exceeding 0.75. Agreement for low-resource languages are notably lower than those for high-resource languages. Low consistency is also pronounced in science reasoning tasks, suggesting these knowledge-intensive problems pose unique challenges for cross-lingual knowledge transfer.
Metric Translation Model Reasoning - Math Reasoning - Science Code generation - Prob. Solving
zh de sw te zh de sw te zh de sw te
spBLEU Gemma2-27B 40.0 51.4 38.2 29.2 80.6 84.8 66.2 57.5 85.5 78.5 76.2 52.3
Llama3.1-70B 35.2 54.4 36.6 35.0 71.8 84.9 64.0 65.3 84.8 78.5 75.5 56.7
Qwen2.5-72B 37.7 50.1 13.5 12.6 77.0 79.4 41.2 40.9 84.6 73.0 48.8 42.1
TER Gemma2-27B 36.2 32.1 40.2 58.6 15.7 12.8 26.9 33.5 15.3 15.6 17.4 33.3
Llama3.1-70B 36.0 30.1 44.0 51.8 19.6 12.9 28.4 26.9 15.1 15.4 17.9 46.8
Qwen2.5-72B 33.1 33.5 76.8 85.7 15.4 16.8 65.8 51.4 14.3 19.6 38.3 53.5
xCOMET Gemma2-27B 86.0 96.1 68.1 71.8 63.2 77.6 36.3 44.1 45.2 46.7 27.0 25.0
Llama3.1-70B 86.8 95.6 66.1 74.0 63.7 77.4 37.1 46.8 43.5 46.3 27.8 28.9
Qwen2.5-72B 87.6 95.6 24.5 30.1 65.2 76.3 20.3 28.4 45.0 45.7 18.0 18.4
Retention Rate Gemma2-27B 1.00 1.08 0.98 0.99 0.98 0.98 1.07 0.81 1.02 0.96 0.89 0.95
Llama3.1-70B 1.01 1.06 1.00 0.97 0.92 1.06 0.99 0.77 1.01 0.98 0.91 0.89
Qwen2.5-72B 1.03 1.04 0.71 0.71 0.90 0.98 1.04 0.79 1.00 1.04 0.93 0.86
Table 9: There exists challenges in domain-specific translation evaluation. The table presents different metric scores of the En-X translation of selected models on specific domains.
5.5 Self-bias is inevitable in the model-based instruction following tasks
Applying model-based evaluation exhibits self-bias, where the judge model prefers the outputs of itself or models from the same family Li et al. (2024); Xu et al. (2024). We further adopt GPT-4o-mini as the judge model, and compute the win-rate against the baseline model GPT-4o. Figure 8 shows that DeepSeek-V3 strongly favors its own outputs, while GPT-4o-mini prefers GPT-4o’s outputs. Nevertheless, the win-rates of other models judged by the different judges are comparable, and the rankings are fairly consistent.
Figure 8: Self-bias is inevitable in model-based instruction following evaluation. DeepSeek-V3 prefers its own outputs, while GPT-4o-mini tends to prefer GPT-4o’s outputs. The win-rates of evaluated models are judged by DeepSeek-V3 and GPT-4o-mini.
5.6 Comaparing open-source and closed-source models on BenchMAX
As demonstrated in Table 10, although GPT-4o-mini and GPT-4o demonstrates strong multilingual capabilities across various tasks, they fall short of DeepSeek-V3, the leading open-source model in our evaluation. This suggests that state-of-the-art open-source models are becoming competitive with their closed-source counterparts. Due to budget constraints, our evaluation of closed-source models is limited to GPT-4o-mini and GPT-4o on some of tasks. A more comprehensive comparison would be valuable for further validating this trend.
Model Rule-based Func Compl. Science
GPT-4o-mini 79.1 78.7 34.1
GPT-4o 80.8 80.1 45.6
DeepSeek-V3 83.9 83.2 47.4
Table 10: The leading open-source model, DeepSeek-V3, bridges the gap to the closed-source models. We compare DeepSeek-V3 and GPT-4o on some of tasks.
6 Conclusion
In this work, we introduce BenchMAX, a comprehensive, high-quality, and parallel multilingual benchmark comprising ten tasks designed to assess crucial capabilities across seventeen diverse languages. The multilingual task data is initially translated from English using machine translation and subsequently refined through multiple iterations of post-editing by native speakers, ensuring high data quality. Through extensive model evaluations, we find that the language-agnostic capabilities of state-of-the-art LLMs remain uneven across different languages. While increasing model size consistently enhances multilingual performance, the performance gap between English and other languages persists, highlighting the need for further efforts to achieve balanced multilingual capabilities.
Appendix A Capability and Task Data Selection
Instruction Following Capability
involves understanding and executing commands accurately and efficiently. In the light of varied evaluation methods - rule-based or model-based - we include two distinct tasks.
  • Rule-based Intruction Following: We collect data from IFEval Zhou et al. (2023), which is a benchmark for evaluating the instruction following abilities of LLMs, composed of around 500 verifiable instructions and can be evaluated for accuracy using automated rules. Note that the accuracy for IFEval is the average of the four accuracies (i.e. prompt-strict, prompt-loose, inst-strict and inst-loose accuracies), following Dubey et al. (2024).
  • Model-based Instruction Following: We collect data from Arena-hard Li et al. (2024) which contains 500 real-world instructions from the Chatbot Arena Chiang et al. (2024), and m-ArenaHard [^6] https://huggingface.co/datasets/CohereForAI/m-ArenaHard which contains translated multilingual versions. This benchmark can provide better model separability and higher alignment with human preference. It is assessed by the Win Rate of the testing model in comparison to the baseline model, GPT-4o, judged against DeepSeek-V3.
Code Generation Capability
refers to automatically producing functional code scripts based on given requirements. Considering variations in difficulty, two separate tasks are included.
  • Function Completion: We collect data from Humaneval+ Liu et al. (2024) which is an augmented version of HumanEval Chen et al. (2021), comprising an expanded test cases. Each problem in the benchmark gives a definition of a Python function accompanied by an English docstring, and requires LLMs to complete the function.
  • Problem Solving: We collect data from LiveCodeBench  [^7] We adopt the code generation subset in LiveCodeBench v4 as the original English dataset. Jain et al. (2024) which provides a more rigorous assessment of the code generation capabilities. It is a much harder benchmark by collecting coding problems in natural language from real competition platforms with live updates.
Long Context Modeling Capability
involves understanding and generating coherent text from extensive input sequences, allowing the model to capture dependencies and relationships within lengthy texts. This paper focuses on the long-context evaluation of multilingual settings based on the RULER benchmark Hsieh et al. (2024).
  • Question Answering: We build synthetic testsets based on RULER, which contains several question answering long-context tasks with pre-defined context length, such as the needle-in-a-haystack (NIAH) test and question answering (QA) test. Since the NIAH test is unrealistic and many models perform perfectly on it, we add a new task called QA-in-a-heystack (QAIAH), where one or several paragraphs are inserted into the haystack. The model then answers the question related to the inserted paragraph instead of finding the obtrusive needle. We reserve the tasks of NIAH, QAIAH, and variable tracking (VT) in our task list, while others are excluded.
Reasoning
encompasses thinking logically, drawing conclusions, making inferences, and solving problems by processing data, applying rules, and utilizing various forms of logic and knowledge representation. Pushing LLMs beyond surface-level tasks, we extend MGSM Shi et al. (2023) and GPQA Rein et al. (2023) requiring deeper understanding and reasoning across different context.
  • Math Reasoning: We collect data from MGSM which evaluates the capability of LLM to solve math reasoning problems in multiple languages, focusing on grade-school level complexity.
  • Science Reasoning: We collect data from GPQA which is crucial for assessing LLM capability for advanced, unsearchable reasoning and critical thinking across diverse, complex domains. It comprises multiple choice questions formulated by experts in the domains of biology, physics, and chemistry, posing extreme challenges where human experts achieve accuracy lower than 70%.
Tool Use Capability
requires the model to translate user queries into executable functions for calling in operating software tools. We extend Nexus Srinivasan et al. (2023) to a multilingual version, which is adopted by Llama3 Dubey et al. (2024).
  • Multiple Functions: Nexus offers a set of functions and user queries. For each query, the language model is required to generate a function call from a list of noisy functions, in accordance with the function definitions and docstrings.
Translation Capability
needs the model to convert text between multiple languages while maintaining semantic meaning accurately. To comprehensively evaluate this capability, we introduce general and task-specific translation datasets.
  • General: General domain data are composed of Flores-200 Costa-jussà et al. (2022), TED Cettolo et al. (2012) and WMT24 Kocmi et al. (2024) testsets. In BenchMAX, we include parallel data from 17 selected languages.
  • Domain: Domain translation data is a by-product of the BenchMAX construction process, encompassing a 17-way parallel task across diverse domains, such as reasoning, code generation, tool usage, and instruction following. Unlike traditional translation tasks, this poses a new challenge to the model by requiring it to determine whether a given segment should be translated or not.
zh es fr de hu ru ja th sw bn te ar ko vi cs sr
w/o special symbols 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68 0.68
+symbol 1 0.91 0.89 0.88 0.92 0.93 0.93 0.91 0.91 0.92 0.95 0.96 0.90 0.88 0.90 0.92 0.99
+symbol 2 0.93 0.93 0.90 0.92 0.95 0.94 0.92 0.93 0.93 0.97 0.99 0.94 0.91 0.92 0.93 1.00
Table 11: The recall of keywords when translating IFEval English data to other languages.
Appendix B Dataset Construction
B.1 Rule-based Instruction Following Dataset
We first filter out some English-specific instructions from the original dataset, such as changing the letter cases. After filtering, the number of remaining samples is 429. The next problem is how to extract the keywords from the translated instruction since the keywords are also translated and are required in the verification step. We try different groups of special symbols to extract translated keywords. The recall rates are presented in Table 5, where order 1 achieves the best performance. Complete results across all languages are provided in Table 11. In addition, the number-word constraints for non-English languages are multiplied by a ratio in order to make the difficulty of the same instruction in different languages comparable. Specifically, we calculate the ratio of the word count of English to that of a non-English language in the Flores-200 corpus using language-specific tokenizers. we also adapt verification rules to multilingual scenarios. For instance, word and sentence segmentation methods may vary across different languages.
During post-editing, we ask human annotators to check whether the translated keywords in the kwargs, which are used by the rule-based program, appear in the translated instruction.
Figure 9: The radar charts visualize the performance of models on each subtask in different languages. Most model evaluated have imbalanced performance across different languages.
B.2 Model-based Instruction Following Dataset
Ten of the sixteen languages required have been provided by m-ArenaHard, which has translated the original dataset into 22 languages using Google Translate. Based on m-ArenaHard, we further translate the English data into six other languages via Google Translate. Subsequently, we ask human annotators to review and edit the translated instructions in all 16 languages.
B.3 Function Completion Dataset
The objective is to translate only the natural texts within the function comments. However, it is challenging to prevent Google Translate from translating other elements, such as function names. Alternatively, we instruct GPT-4o to complete this translation task with well-designed prompts (Table 18). Furthermore, a human post-editing process is employed to refine the quality of the generated translation.
B.4 Problem Solving Dataset
Similar to the Function Completion Dataset, we employ GPT-4o to translate the English problems into other 16 languages with a well-designed prompt (Table 19), since Google Translate cannot distinguish the parts that should remain untranslated. Human review is also used to ensure the overall quality of the translated texts.
B.5 Math Reasoning Dataset
Given that the MGSM examples are written in ten languages we need, we only translate the English version into the remaining six languages via Google Translate. This is also followed by a manual checking procedure.
B.6 Science Reasoning Dataset
The question and the four options of each sample are translated into 16 other languages by Google Translate. In particular, the question and options are concatenated by option markers like “(A)”. After translation, we extract the translated question and options to form a new sample.
B.7 Long-Context Question Answering Dataset
The haystacks, needles, paragraphs and questions related to QAs are translated to other languages. We use the parallel testsets from the UN corpus Ziemski et al. (2016) as the haystack. The English version contains about 128k tokens, and we extend it to other languages using Google Translate. The sentence of the needle is also translated into 16 other languages, in which UUIDs are employed as keys and values that are not translated. With respect to the QA data, we translate the paragraphs and questions in XQuAD Artetxe et al. (2020) to the languages we need. Note that we also use the trick in translating IFEval to extract the answer spans. With access to our multilingual haystacks, needles and paragraphs, we are able to synthesize the multilingual long-context testsets.
B.8 Multiple Functions Dataset
We only translate the user queries from English into other languages, given that the majority tool descriptions are written in English. The user queries are initially translated by Google Translate and subsequently adjusted by human annotators. To preserve the English parameters, we replace them with placeholders before machine translation and restore them afterward.
Appendix C Model Information
Here we list the evaluated models in this section.
Llama3.1-Instruct Dubey et al. (2024)
series contains three multilingual large language models with number of parameters ranging from 8B to 405B. The pre-training corpus of Llama3.1 contains 8% multilingual tokens, and multilingual alignment is also optimized during post-training. In our experiments, we evaluate the 8B version and the 70B version of Llama3.1-Instruct.
Qwen2.5-Instruct Qwen Team (2024)
is a collection of multilingual language models with several sizes, ranging from 0.5B to 72B. The models are trained with multilingual tokens in both pre-training stage and post-training stage, and are rigorously evaluated on several multilingual tasks. In our experiments, we evaluate the 7B, 32B ,and 72B version of Qwen2.5-Instruct.
Aya-Expanse Dang et al. (2024)
is an open-weight research of models with advanced multilingual capabilities, supporting 23 languages. The Aya Expanse 8B and 32B variants are instruction-tuned and beat Llama3.1-instruct models on the m-ArenaHard, a multilingual instruction following benchmark.
Gemma2-IT Team et al. (2024)
family demonstrates strong multilingual capabilities, although this is not highlighted in the technical report. We benchmark the 9B and 27B variants of Gemma2-IT.
InternLM2.5-chat Cai et al. (2024)
is the successor of InternLM Team (2023), which is claimed as a multilingual model. We include the 7B version and 20B version in our experiments. InternLM2.5-7B-chat-1m is a long-context variant supporting context windows with 1M tokens.
DeepSeek-V3 DeepSeek-AI et al. (2024)
is one of state-of-the-art open-source models that achieve performance comparable to that of the best proprietary models. It is a 671B MoE model, with 37B activated for each token. A multilingual corpus and a multilingual-optimized tokenizer are incorporated into their training process.
DeepSeek-R1-Distill-Llama & DeepSeek-R1-Distill-Qwen
Guo et al. (2025) are dense models with long reasoning capabilities, and are distilled from DeepSeek-R1 based on Llama3.1-8B, Llama3.3-70B-Instruct, Qwen2.5-Math-7B, and Qwen2.5-32B.
GPT-4o & GPT-4o-mini OpenAI (2024)
are two of the best proprietary models that also achieve remarkable performance on multilingual tasks. Their tokenizer can better compress multilingual texts than that of GPT-4. GPT-4o-mini is the smaller version of GPT-4o with powerful performance.
Appendix D Details about Prompt Templates
We present the prompt templates used in each task in this section. Table 12 and Table 13 show the native-CoT prompts for MGSM and GPQA. Table 14 shows the prompt templates for some tasks where the original English template is used. Table 15 shows the prompt templates of the long-context modelling task. Table 17 shows the LLM-Judge Instruction for comparing two translations.
Table 12: The native-CoT prompts of the mathematical reasoning task.
Table 13: The native-CoT prompts of the scientific reasoning task.
Task Prompt Template
Rule-based instruction following {prompt}
Model-based instruction following {prompt}
Problem Solving [System Message]
You are an expert Python programmer. You will be given a question (problem specification) and will generate a correct Python program that matches the specification and passes all tests. You will NOT return anything except for the program.
[User Message]
### Question:
{question}
### Format: Read the inputs from stdin solve the problem and write the answer to stdout (do not directly test on the sample inputs). Enclose your code within delimiters as follows.
```python
# YOUR CODE HERE
### Answer: (use the provided format with backticks)
Function Completion [User Message]
Please provide a self-contained Python script that solves the following problem in a markdown code block:
{prompt}
[Assistant Message]
Below is a Python script with a self-contained function that solves the problem and passes corresponding tests:
```python
Tool use [Tool Info]
{prompt}
Table 14: The prompt templates of the listed tasks. The prompt in the template is multilingual.
Subtask Prompt Template
NIAH [User Message]
Some special magic uuids are hidden within the following text. Make sure to memorize it. I will quiz you about the uuids afterwards.
{heystack}
What are all the special magic uuids for {query} mentioned in the provided text?
[Assistant Message]
The special magic uuids for {query} mentioned in the provided text are
QA in a heystack (QAIAH) [User Message]
Answer the questions based on the given documents. Only give me the answers and do not output any other words.
The following are given documents.
{context}
Answer the questions based on the given documents. Only give me the answers and do not output any other words.
Questions:
{query} [Assistant Message]
Answers:
Variable Tracking (VT) [User Message]
Memorize and track the chain(s) of variable assignment hidden in the following text.
{context}
Question: Find all variables that are assigned the value {query} in the text above.
[Assistant Message]
Answer: According to the chain(s) of variable assignment in the text above, 5 variables are assgined the value {query}, they are:
QA [User Message]
Answer the question based on the given documents. Only give me the answer and do not output any other words.
The following are given documents.
{context}
Answer the question based on the given documents. Only give me the answer and do not output any other words.
Question: {query}
[Assistant Message]
Answer:
Table 15: The prompt templates of the long-context modelling task.
Score the following translation from {src_lang} to {tgt_lang} with respect to the human reference on a continuous scale from 0 to 100 that starts with “No meaning preserved”, goes through “Some meaning preserved”, then “Most meaning preserved and few grammar mistakes”, up to “Perfect meaning and grammar”
{src_lang} source: “{source}”
{tgt_lang} translation: “{target}”
Score:
Table 16: The GEMBA-SQM prompt.
[System Message]
Please act as an impartial judge and evaluate the quality of the lang translations provided by two humans for the English source sentence displayed below. You will be given human A’s translation and human B’s translation. Your job is to evaluate which human’s translation is better.
You must identify and correct any mistakes or inaccurate information.
Consider if the human’s translations are accurate and fluent. Accurate means the translation conveys the same meaning, information, and nuances as the original source text. Fluent refers to the quality of the translation in terms of its naturalness, readability, and adherence to the grammatical, stylistic, and idiomatic conventions of the target language.
Then consider whether the human’s translations are consistent with the context. Code input/output and programming language syntax should not be translated. Finally, review the formatting of the translated text, including indentation, to ensure it is consistent and appropriate.
After providing your explanation, you must output only one of the following choices as your final verdict with a label:
1. Human A is significantly better: [[A>>much-greater-than$>$ >B]]
2. Human A is slightly better: [[A$>$B]]
3. Tie, relatively the same: [[A$=$B]]
4. Human B is slightly better: [[B$>$A]]
5. Human B is significantly better: [[B>>much-greater-than$>$ >A]]
Example output: “My final verdict is tie: [[A$=$B]]”.
[User Message]
<|||| Source Text||||>
{source}
<||||The Start of Human A’s Translation||||>
{translation_1}
<||||The End of Human A’s Translation||||>
<||||The Start of Human B’s Translation||||>
{translation_2}
<||||The End of Human B’s Translation||||>
Table 17: LLM-Judge Instruction
[System Message]
You are a professional translator specializing in technical content. Please translate the following English Python codes into {tgt_lang}, adhering to these specific guidelines:
1. **Do not translate** content representing code input/output or programming language syntax. Only translate content in comments.
2. **Maintain the original formatting** of the text, structure and indentation.
3. **Do not translate** any LaTeX code.
4. **Only output the translation** without any additional comments or explanations.
[User Message]
{problem}
Table 18: Prompt for translating the Function Completion task.
[System Message]
You are a professional translator specializing in technical content. Please translate the following English coding problems into {tgt_lang}, adhering to these specific guidelines:
1. **Do not translate** any LaTeX code.
2. **Do not translate** content representing code input/output or programming language syntax.
3. **Maintain the original formatting** of the text and structure.
4. **Only output the translation** without any additional comments or explanations.
[User Message]
{problem}
Table 19: Prompt for translating the Problem Solving task.
Appendix E Detailed results
Figure 9 illustrates the detailed results of each model on each task.